If a diffraction grating produces a third-order bright spot for red light (of wavelength 700 nm) at 65.0° from the central maximum, at what angle will the second-order bright spot be for violet light (of wavelength 400 nm)?

1
Expert's answer
2022-01-13T09:27:28-0500

"\theta _3=65.0°"

"\lambda_1 =700nm =700×10^{-9}m"

"\lambda _2=400nm=400×10^{-9}m"

We know that angular position of a bright band is given by

"Sin\theta_m=\frac{m\lambda}{d}"

For the third-order bright band,in the first wavelength "\lambda _1" whereas "m" "=+-3"


"Sin\theta _3=\frac{3\lambda_1}{d}"


"d=\frac{3\lambda_1}{Sin\theta_3}"


"d=\frac{3×700×10^{-9}}{Sin65.0°}"


"d=2.32\u00d710^{-6}m"


"Sin\theta_2=\frac{2\lambda_2}{d}"

Hence

"\theta_2=Sin^{-1}[\frac{2\lambda2}{d}]"


"\theta_2=Sin^{-1}[\frac{2×400×10^{-9}}{2.32×10^{-6}}]"


"\theta_2=20.2°"